Can you please help me with this problem If it is not too much to ask can you please walk me through each step so that I can complete the actual assigned problems. Thank you so much!!!!
How many ounces of pure gold must be added to five ounces of something that is 1% gold in order that the mixture is 6% gold?
Thanks again for everything!!!

Answer by scott8148(5891)

let x="ounces to be added"

5(1%)+x(100%)=(5+x)(6%) ___ .05+x=.30+.06x ___ .94x=.25

the amounts multiplied by the concentrations (%) have to balance